### v1

# Insight-HXMT View of the BHC Swift J1727.8-1613 during its outburst in 2023

Abstract
The transient Galactic black hole candidate Swift J1727.8-1613 went through an outburst for the very first time that started in August 2023 and lasted for almost 6 months. We study the timing and spectral properties of this source using publicly available archival Insight-HXMT data for the first 10 observation IDs that last from MJD 60181 to 60198 with a total of 92 exposures for all three energy bands. We extracted the quasi-periodic oscillation properties by model fitting the power density spectrum and from those properties we designate that the QPOs are type-C in nature. We also conclude that the origin of the QPOs could be the shock instabilities, formed in the transonic accretion flow around black holes. The spectral analysis was performed using simultaneous data from the three on-board instruments LE, ME, and HE of HXMT in the broad energy band of $2-150$ keV. To achieve the best fit, we needed a combination of interstellar absorption, power-law, multi-color disk-blackbody continuum, gaussian emission/absorption, and power-law reflection by neutral material. From the spectral properties, we found that the source was in an intermediate state at the start of the analysis period and was making a transition toward the softer states. The shock (the boundary layer of the corona) moved inward in progressive days in accordance with the spectral nature. We found that the source is present in a high-inclination binary system. The hydrogen column density was found with an average value of $0.27_{-0.17}^{+0.08}\times 10^{22}$ cm-2.
X-rays: binaries – stars: black holes – stars: individual (Swift J1727.8-1613) – accretion, accretion disk – shock – radiation
journal: ApJ
1 Introduction
Black holes (BHs) are one of the end-products of massive stars. Stellar-mass black holes (SBHs) are one of the three types of presently detected BHs. They reside in binary systems with a companion star and most of their radiation comes out in the form of X-rays. Thus, they are called black hole X-ray binaries or BHXRBs. Depending on the mass of the companion star, they can be mainly of two types: low-mass X-ray binaries (LMXRBs) or high-mass X-ray binaries (HMXRBs). LMXRBs consist of type-A or later type of stars whereas the HMXRBs consist of giant O or B-type of stars (White et al. 1995; Tetarenko et al. 2016). Matter coming from the companion accumulates at a distance from the BH due to lack of viscosity, known as the pile-up radius (Chakrabarti et al. 2019; Chatterjee et al. 2022). Due to the accumulation of more matter, the temperature will rise. This gradual increment of temperature sets in instability at the pile-up, which therefore increases the viscosity. At some point in time when the accumulated matter will gain enough viscosity to push forward, accretion starts. On the other hand, irradiation of the accumulated matter at the outer boundary by the central object can also trigger mass accretion when the temperature of the accumulated matter crosses a critical limit (King & Ritter 1998; Mondal 2020). Thus, these systems make ideal laboratories for studying the physics of accretion. The BHXRBs mostly stay in a quiet dormant state when their flux stays at a very low detection level. Episodically, the radiation level increases by a very large factor and they become easily detectable. This is known as an outburst (Remillard & McClintock 2006). The X-ray luminosity increases by several orders of magnitude during an outburst when compared to a quiescence. Depending on the nature of the outburst, the sources are classified into two types: persistent and transient sources. Most of the BHXRBs are transient type of sources. They stay in the dormant state for most of their lifetimes. The luminosity becomes $\sim 10^{35}$ erg s-1 during episodic outbursting phase.
During the onset of an outburst, the flux in the light curve changes noticeably. According to nature, the outbursts are divided into two types (Debnath et al. 2010): fast rise slow decay (FRSD) and slow rise slow decay (SRSD). However, there are also some other classified outburst types, e.g., according to Zhang et al. (2019), outbursts can also be classified into glitch, reflare, multipeak, and mini-outburst. There are generally four designated spectral states during a complete BH outburst (Remillard & McClintock 2006). They are known as hard state (HS), hard intermediate state (HIMS), soft intermediate state (SIMS), and soft state (SS). Generally, when an outburst starts, it starts in the HS. Then it slowly moves toward the SS through the HIMS and SIMS (Belloni et al. 2005, 2011, and references within). After it reaches its SS the rising phase generally comes to an end and the outburst starts its decaying phase. In the decaying phase, it goes in the opposite direction and reaches the HS again. When a BH outburst goes through all these states, it is known as a type-I outburst. When the soft state is absent, it is known as a ‘failed’ or type-II outburst.
The radiation spectrum of a BH generally consists of two main components. One is the soft thermal multi-colour blackbody component, which is due to the blackbody radiation of the seed photons in the standard accretion disk (Shakura & Sunyaev 1973; Novikov & Thorne 1973). The other is the hard power-law component, which extends up to very high energies. The origin of this component is due to the inverse Comptonization of a fraction of seed photons from the disk, which are intercepted by a hot Compton cloud (Sunyaev & Titarchuk 1980, 1985; Haardt & Maraschi 1993; Zdziarski et al. 1993; Titarchuk 1994; Chakrabarti & Titarchuk 1995; Życki et al. 1999). When the outburst is in HS, there is very small contribution from the thermal component and mostly the non-thermal component dominates. In SS, the thermal component becomes dominant while in the intermediate states (HIMS and SIMS), the contribution from the two components stays comparable. other than these two components, other components could also be present in the radiation spectrum of a BH. In the case of a high soft state (HSS), there could be the presence of bulk motion Comptonization (BMC) which takes place due to the relativistic speed of matter when it reaches a very close distance to the BH (Blandford & Payne 1981a,b; Payne & Blandford 1981; Psaltis & Lamb 1997; Borozdin et al. 1999; Psaltis 2001;). It was been recently suggested by Chatterjee et al. (2023) that BMC was present in the radiation spectrum of the BHC MAXI J0637-430. The reprocessed emission plays an important role in the case of BHXRBs. Fraction of electromagnetic radiation can come back from the optically thin and geometrically thick Compton cloud to the optically thick and geometrically thin accretion disk due to the strong gravitational bending effect of the BH (Riaz et al. 2021). The illumination of the Comptonized photons from the Compton cloud to the accretion disk is known as returning or reprocessed radiation (George & Fabian 1991; Ross & Fabian 2005; Garcia & Kallman 2010). The reflection spectrum is a very important part of the whole radiation spectrum which helps determine many important properties of the source. The reflection radiation can produce some very specific features in the spectrum. One of which is the iron $K_{\alpha}$ emission line at $\sim 6.4$ keV. This happens when the returning radiation illuminates the disk and there is the fluorescence of ionized Fe. This line emission has a width of the order of $\sim 1~{}eV$ (Pozdniakov et al. 1979). However, for rotating black holes there are effects like Doppler shift and gravitational broadening which makes the line broadened and for rapidly rotating black holes this line can have two peaks due to red-shifting and blue-shifting of the component. This line emission is a very important aspect in estimating the spin of the BH. There could also be other features like Fe-K absorption edge and Compton hump. These features arise because of counterplay between two phenomena. Which feature will be prominent will be resulted in depending on which phenomenon is dominant. These are also energy-dependent. The interplay between photoelectric absorption and Compton scattering phenomena (Guilbert & Rees 1988; Lightman & White 1988; Fabian et al. 1989) will decide the spectral feature. At higher energies above $20$ keV, the scattering will be dominant resulting in a Compton hump, whereas, in low energies, there will be the dominance of photoelectric absorption and the spectrum drops away (Lightman & White 1988). As these features take place at a closer distance from the compact object (here a BH), we can notice broad and asymmetric reflection profiles in the spectrum due to distortion due to high gravitation. Proper modeling of the composite spectrum using different combinations of models can give good information about the physics that is going on there. In this regard, high signal-to-noise (s/n) data in the broadband range, e.g., Insight-HXMT (Zhang et al. 2020) can be very useful.
Apart from the spectral features and variabilities, there are temporal variabilities also. Time variability, which is proportional to the length scale of the photon-emitting region, is as important as X-ray spectroscopy to understand the physical properties of accretion flow dynamics in the region near to the BH. During an outburst, we notice that there are very small time-scale variabilities in the light curves, especially in the high-energy ones. If we make a Fourier transformation of the light curve, we notice that fast variability is present in the power density spectrum or PDS (van der Klis 1989). In the PDS, which is a variation of $P(\nu)$ with $\nu$, the variability is mainly aperiodic, which is a direct consequence that the accretion flow onto the Bh is turbulent itself. The aperiodic variability, which is the result of the combination of the variation of the source flux and the variation in Poisson statistics, contains both broadband noise and narrow features. The broadband noise is spread over a large frequency range, whereas the narrow feature is a power peak in narrow frequency ranges. The narrow peak is known as the quasi-periodic oscillation (QPO), which is one of the most important properties of a BH. There is a break frequency ($\nu_{b}$) that also characterizes the PDS (van der Klis 1994). The noise is modelled by the power-law function, whereas the QPOs are modelled by one or multiple Lorentzian profiles. In BHs, we see low-frequency QPOs (LFQPOs) mainly due to their geometrical origin. Depending on some properties like the frequency ($\nu$), $Q$-value ($=\nu/\delta\nu$, where $\delta\nu$ is the full width at half maximum or FWHM), RMS, etc., LFQPOs are classified into three different types: type A, B, and C (Casella et al. 2005). Although the clear origin of QPOs is still a matter of debate at present, several ideas were put forward as to why these features are being observed. Explanations like Lense-Thirring precession (Stella et al. 1999; Ingram et al. 2009), magneto-acoustic waves (Titarchuk et al. 1998), accretion-ejection instability (Tagger & Pellat 1999), shock oscillation model (Molteni et al. 1996; Chakrabarti et al. 2005, 2008, 2015) were put forward to explain the origin of LFQPOs in stellar-mass BHs. The shock oscillation model can give a simultaneous explanation of both the spectral and temporal properties, which makes it a very generally comprehensive model. High-resolution light curve data becomes very useful to study the temporal variability properties of BH. In this regard, Insight-HXMT is also very useful as well as studying the spectral properties.
Apart from these properties and features of a BH outburst, there are a few other estimations that give a direct correspondence between the spectral and temporal properties. A good correlation is observed between the two using just the light curve properties, e.g., the hardness ratio or HR, hardness intensity diagram or HID (Homan et al. 2001). Some model-fitted approaches are also important in this regard, e.g., accretion rate ratio or ARR, accretion rate ratio intensity diagram or ARRID (Mondal et al. 2014; Jana et al. 2016; Chatterjee et al. 2020). The RMS-intensity diagram or RID (Munoz-Darias et al. 2011), and hardness ration-intensity diagram or HRD (Belloni et al. 2005) are also very important estimates when it comes to correspondence between spectral and temporal properties. In different spectral states, all of these properties show different values, which are quite general for BH outbursts.
The BHC Swift J1727.8-1613 was discovered very recently on 24 August 2023 or MJD 60180 (Kennea & Swift Team 2023). The source was very bright as MAXI (Matsuoka et al. 2009) detected a flux of $7~{}Crab$ in the $2-20$ keV energy range. After that multi-wavelength follow-up observations were carried out by several ground-based and space-bourne telescope facilities (Baglio et al. 2023; Miller-Jones et al. 2023; Negoro et al. 2023b; O’Connor et al. 2023; Wang & Bellm 2023; Williams-Baldwin et al. 2023). The source is reported to have a distance of $2.7\pm 0.3kpc$ (MataSanchez et al. 2024) and a mass of $10\pm 2~{}M_{\odot}$ (Svoboda et al. 2024). The revelation of the presence of hydrogen and helium emission lines in the optical spectrum lead to its classification as an LMXRB (Castro-Tirado et al. 2023). The X-ray spectrum and other properties point to its nature as a BHXRB (Liu et al. 2023; Sunyaev et al. 2023). This was supported strongly after the detection of type C QPOs (Palmer & Parsotan 2023; Draghis et al. 2023; Bollemeijer et al. 2023; Mereminskiy et al. 2023) and flat radio spectrum (Miller-Jones et al. 2023; Bright et al. 2023). There is X-ray monitoring of the source with many satellites, e.g., NICER, NuSTAR, Insight-HXMT, IXPE, etc. Using IXPE data on September 7, 2023, Veledina et al. (2023) estimated its polarization with a polarization degree (PD) of $4.1\%\pm 0.2\%$ and PA of $32^{\circ}.2\pm 1^{\circ}.$3. There was a further polarization study of the source by Ingram et al. (2023). The presence of evolving LFQPOs at the high energy band of using AstroSat data was reported by Nandi et al. (2024). Using X-ray spectroscopy, there were reports about the spectral nature of the source. At the onset of the outburst, the spectral nature was consistent with a hard state with a photon index of power-law ($\Gamma$) $\sim 1.5-1.7$ (Liu et al. 2023). Then using the NICER observations, Bollemeijer et al. (2023) reported the softening of the spectrum with a substantial contribution of the soft blackbody component after 25 days of the start of the outburst. The presence of LFQPOs was reported by Draghis et al. (2023). Using NICER data on August 29 2023 (MJD 60185), Debnath et al. (2023) reported the presence of a soft time lag of $0.014\pm 0.001$ sec between energy bands of $3-10$ and $0.5-3$ keV at QPO frequency. With reflection spectroscopy, they found the inclination to be $\sim 87^{\circ}$.
Here is how this paper unfolds. In §1, we give an introduction. In §2, we describe the observation, data selection, reduction, and analysis. In §3, we portray the results. In §4, we make the discussion about the possible physical origin of our result. And finally, in §5, we make a summary of our results and draw some conclusions.
2 Observation, Data Selection, Reduction, and Analysis
This source is being extensively monitored by various X-ray satellites, e.g., NICER, NuSTAR, Insight-HXMT, IXPE, etc. For our analysis, we use China’s first dedicated X-ray satellite Insight-High X-ray Modulation Telescope (HXMT; Zhang et al. 2020). In the subsequent subsections, we discuss the data selection, reduction, and analysis, respectively.
2.1 Data Selection
This source was regularly monitored by the Insight-HXMT satellite. In the data archive, observations are available on-demand. At the time of the work, we found that there was a total of 34 observation IDs were available, although 3 of them were not downloadable. Out of those usable 31 observation IDs, we took the first 10 observation IDs. However, unlike other X-ray satellite data, each observation ID in HXMT has several different exposures. Also, some of the observation IDs continue for almost 2 to 3 days. We found that for these first 10 observation IDs, each of them has a minimum of 7 to a maximum of 18 exposures. We checked the analysis of all the exposures for the first 3 observation IDs and found that no significant change was noticed. Thus we did not fit every exposure for each observation ID. The details about the selected data are given in Table 1.
Table 1: List of Data used
Obs. Id.[1] Start Date[2] End Date[2] Start MJD[3] Stop MJD[3] Exp. (s)[4]
(1) (2) (3) (4) (5) (6)
P0614338001 2023-08-25 2023-08-27 60181.3410 60183.0597 148495
P0614338002 2023-08-27 2023-08-28 60183.0597 60184.2327 137343
P0614338003 2023-08-29 2023-08-31 60185.3075 60187.1589 159958
P0614338004 2023-08-31 2023-09-02 60187.1588 60189.0765 165684
P0614338005 2023-09-02 2023-09-04 60189.0765 60191.0601 171384
P0614338006 2023-09-04 2023-09-06 60191.0601 60193.4447 206034
P0614338008 2023-09-07 2023-09-08 60194.0350 60195.0926 91378
P0614338009 2023-09-08 2023-09-09 60195.0926 60196.0840 85659
P0614338010 2023-09-09 2023-09-10 60196.0840 60197.0754 85653
P0614338011 2023-09-10 2023-09-11 60197.0753 60198.0667 85650
[1] Column 1 lists each observation ID.
[2] In columns 2 and 3, we give the start and end date of each observation ID.
[3] In columns 4 and 5, we give the start and end MJD of each observation ID.
[4] Column 6 gives the exposure time of each observation ID.
2.2 Data Reduction
Insight-HXMT, China’s first astronomical X-ray satellite, was launched on 2017 June 15 (Zhang et al. 2014, 2020) to observe and study the properties of black holes, neutron stars (NSs), active galactic nuclei (AGNs), and other astronomical phenomena. The main scientific instrument is an array of 18 NaI/CSI phoswich scintillation detectors, each of which covers an effective area of 286 cm2. The satellite has three instruments or payloads: High Energy (HE; Liu et al. 2020), Medium Energy (ME; Chen et al. 2020), and Low Energy (LE; Cao et al. 2020). They have an effective area of 5100, 952, and 384 cm2 and cover an energy range of $20-250$, $5-30$, and $1-15$ keV, respectively.
After downloading on-demand level-1 data from the archive, we produced cleaned level-2 data for science analysis. This process of cleaning raw data was done in the following way. We first installed the HXMTDAS [^1] http://hxmt.org/index.php/usersp/dataan v2.05. Using this software, we run the hpipeline command using proper input and output directories. This pipeline performs several subsequent automatic commands for all three instruments. However, some suitable conditions are needed to be provided. Some conditions were set to get a good time interval, e.g., elevation angle $>10^{\circ}$; geomagnetic cutoff rigidity $>8~{}GeV$; pointing offset angle $<0.04^{\circ}$; $>600~{}s$ away from the South Atlantic Anomaly (SAA). All these commands altogether extract and clean the raw data and produce cleaned and analyzable science products. This is discussed in the HXMT manual [^2] http://hxmten.ihep.ac.cn/SoftDoc/501.jhtml in full details. The specific commands hespecgen, mespecgen, and lespecgen produce the spectra for HE. ME, and LE instruments. Whereas, the helcgen, melcgen, and lelcgen commands produce the light curve files for the three instruments. herspgen, merspgen, and lerspgen produce suitable response files. The background subtraction for both the timing and spectral data was performed using the commands hebkgmap, mebkgmap, and lebkgmap for instruments HE, ME, and LE, respectively. For $\chi^{2}$ fit-statistics in XSPEC, we group the spectrum using the grppha task of FTOOLS to a minimum of 30 counts per bin. We also set the time bin size to $0.01s$ to produce suitable light curves for generating PDS and looking for QPOs.
Using the cleaned spectra and light curves, we analyze the source properties, which are discussed in detail in the very next subsection.
2.3 Data Analysis
We perform timing and spectral analysis of the very first outburst during 2023 of the BHC Swift J1727.8-1613. We have downloaded the daily average light curve using the MAXI/GSC [^3] http://maxi.riken.jp/star_data/J1727-162/J1727-162.html public archive. Using light curves from all three modules of the HXMT, i.e., LE, ME, and HE, we have produced $0.01~{}s$ time-binned light curves first. Using those light curves, we have produced power density spectrum (PDS) by using the fast Fourier transformation (FFT) in the powspec task of the XRONOS package in the HEASoft software. Each observation’s data was subdivided into several intervals, each of which contains 8192 newbins. First, the PDS for each interval is generated and then they are averaged to make a resultant PDS. The PDS is normalized in a way that their integral will give the rms-squared fractional variability. White noise level is subtracted and a geometrical rebinning of -1.02 is used. For some data, we needed to use a geometrical rebinning of -1.05 to have a prominent QPO-like nature in the PDS. Using these steps, we searched for LFQPOs. We fit the QPO shape by the use of a Lorentzian model in powspec and extract QPO properties like frequency ($\nu_{qpo}$), full-width at half maximum or FWHM, normalization. For many observations, we have found the presence of single or multiple harmonics. We have also extracted their properties by fitting them with a Lorentzian model. We fitted the light curves of all the exposures of the listed 10 observations. We report them in the result section.
For spectral analysis, we have also used all three modules, i.e., LE, ME, and HE and fitted the broadband data in the $2-150$ keV energy band. Looking for the best fit, we have found that using the combination of disk blackbody, power-law, gaussian, and pexrav models, the data is best fitted. For some observations, we needed to add a gabs model to achieve the best fit. For interstellar absorption, we have used the tbabs model. Since we are simultaneously fitting all three modules, we have included a constant to normalize the three resultant fittings. The following two are our model combination: i) constant*tbabs*gabs*(diskbb + power-law + gaussian + pexrav) (for those for which gabs was required), ii) constant*tbabs*(diskbb + power-law + gaussian + pexrav) (for those for which gabs was not required. The spectral analysis was not done for all the exposures of the listed observations. Like the timing analysis, we did not spectrally analyze all the exposures, given in Table A. The spectrally analyzed exposures are marked with ‘*’ in Table A. The reason behind this is that, for timing analysis, we found variations of timing properties within a single day, whereas, for spectral analysis, the properties do not vary that much within such a short time. We report them in the result section.
We note that light curves for the LE instrument were not produced for some exposure IDs although we run the same pipeline for all three instruments. Thus we only represented those exposures MJDs in Table 3 for which all three light curves were available for analysis for uniformity.
3 Results
Using the timing and spectral analysis, we have studied the accretion flow properties of the very recent 2023 outburst of the BHC Swift J1727.8-1613. We studied the QPO properties of the source during the outburst as well as the spectral nature and radiation properties of the source using Insight-HXMT data. We broadly discuss the results in the two subsections below.
3.1 Temporal Properties
Here, we discuss the evolution of the flux and hardness ratio, and QPO properties during the first 10 observation IDs of the outburst.
Figure 1: Variation of the (a) MAXI/GSC $2-20$ keV count rate, (b) MAXI/GSC $2-6$ and $6-20$ keV count rates, and (c) hardness ratio with time. The HR is the ratio of the $6-20$ keV count rate to the $2-6$ keV count rate of the MAXI/GSC data.
3.1.1 Outburst Profile and Hardness Ratio
In Figure 1, we show the variation of the flux during the outburst. The outburst started roughly at around MJD 60180 (2023 August 24), when its flux came out of quiescence. The MAXI/GSC flux started to rise after this date, as we can notice from panel (a) of Figure 1. Within five days, the flux increased very rapidly and the total flux reached its peak on MJD 60185 (2023 August 29). Then the flux started to decrease very slowly. It decreased very slowly and gradually except for the period from MJD 60200 to 60225, where the flux showed a significant variation. In panel (b), we show the variations of the $2-6$ and $6-20$ keV energy bands count rates with time. We notice that, after MJD 60185, the hard flux (i.e., $6-20$ keV) decreased very slowly and gradually without showing any other variation, whereas, the soft $2-6$ keV flux showed a lot of variation just like the total $2-20$ keV flux. MJD 60185 is the peak for both the total $2-20$ keV and hard $6-20$ keV fluxes, whereas, the soft flux reached its peak much later at around MJD 60217. This suggests that the variation in the total count rate is mainly due to the variation of the soft component. After this period, the soft flux also decreased slowly to reach the end of the outburst and enter again into the quiescence phase.
In the panel (c), we show the variation of the hardness ratio, which is the ratio of the hard $6-20$ keV to the soft $2-6$ keV flux. At the start of the outburst, the HR was $\sim 0.5$. From this value, it first gradually decreased to MJD 60235 as the outburst progressed. After that, it became almost constant, although decreasing very slowly to $\sim 0.025$ at the end of our taken data of the outburst on MJD 60350. The hardness ratio is defined as $F_{H}/F_{S}$, where $F_{H}$ and $F_{S}$ are the hard and soft fluxes respectively. When this value is on the higher side, the spectral nature of the source is hard. This is the general case when an outburst is starting. When this value is low, there is dominance of soft flux over hard flux and the spectral nature would be soft. Without studying the detailed spectral analysis, HR provides a quick and rough idea about the spectral nature of an outburst. However, for this source, we notice that HR was a little higher at the onset of the outburst, where the hard flux was also comparable to the soft flux. However, after a few days, soft flux completely took over the hard flux, and the HR was also not high. Depending on the HR value, we propose that the source was already in the SIMS after the onset of the outburst. After MJD 60185, it made a transition into the SS. Since we are not analyzing the full duration of the outburst, it is not the scope of the paper to study the full spectral classification. We only spectrally define the source in the duration which falls into our analysis range.
3.1.2 Quasi Periodic Oscillations (QPOs)
As described in §2, we used the $0.01s$ time-binned light curves from all three bands (LE, ME, and HE) for our timing analysis. Using those light curves, we produced the power density spectrum (PDS) of those light curves to study QPOs. We find that QPO nature was present in all of the light curves in all three energy bands. In Figure 2, we show the variation of the QPO frequency ($\nu_{qpo}$) for the 10 different obs. IDs that we have used for our timing analysis. We first checked the first obs. ID P0614338001. It has several exposure IDs (given in Table 2). While we fit every exposure ID of this obs ID, we found that the QPO frequency was changing even within this small period. This is why we fitted all the light curve data of the exposures of the obs IDs that we have taken for analysis. We find that for every single obs ID, the $\nu_{qpo}$ showed significant variation. Figure 2(a), (b), (c) show the variations of the $\nu_{qpo}$ with time for 10 different obs IDs for LE, ME, and HE respectively.
Evolution of QPO frequency in different observation ID.s
Refer to captionRefer to caption
Figure 2: Evolution of QPO frequencies (Hz) with time. Here we have shown three different plots for (a) LE, (b) ME, and (c) HE instruments. Panels 1–10 in each of the three of these plots (a, b, c) represent the $\nu_{qpo}$ variation for the 10 observation IDs we used.
Figure 3: Evolution of QPO frequency with time during the whole period of analysis for (a) LE, (b) ME, and (c) HE.
In Figure 3, we show the evolution of the $\nu_{qpo}$ for the total duration of our analysis period. In (a), (b), and (c) panels the variations are for LE, ME, and HE respectively. We find that the QPO frequencies were $\sim 0.23\pm 0.008$, $0.24\pm 0.002$, and $0.24\pm 0.002$ Hz for LE, ME, and HE on the starting day of our analysis. After this day, the $\nu_{qpo}$ shows a rapid increase with time till $\sim$ MJD 60187 when $\nu_{qpo}$ was $\sim 1.2$Hz for all the three bands. After this date, it increased very slowly till MJD $\sim 60190$ when $\nu_{qpo}$ was $\sim 1.45$. Then it decreased for a very short period until it very rapidly increased to $1.5$Hz again in all three bands. After that, it showed varying nature till MJD $\sim 60197$ and then it finally increased again. This might be caused by the varying nature of the origin of these oscillations. We will discuss this in the next section.
Using the PDS fitting, we extracted some information (e.g., full-width at half maximum (FWHM), Normalization (LN)) of QPOs. We also extracted source and background count rates from the exposures. Using these estimates, we calculated the $Q$-value and RMS that represent the sharpness of the QPO and the fractional variability in the PDS, respectively. The values are listed in Table 3 in columns 5–7 ($Q$-value) and 8–10 (RMS) for LE, ME, and HE respectively.
Evolution of rms (%) with QPO frequency and time
Refer to captionRefer to caption
Figure 4: 3-D representation of the variation of $\nu_{qpo}$ and (%)RMS with time (MJD) for (a) LE, (b) ME, and (c) HE.
In Figure 4, we show a 3-D representation of the evolution of QPO RMS with corresponding frequency. In the LE band (Fig.4a), The RMS was higher at the start of the analysis period and it then gradually decreased. In the ME band (Fig. 4b), it increased from the onset of the analysis period for some time and then again decreased. HE band (Fig. 4c) shows a similar trend as of LE band. Since the QPO nature has already been reported in the recent paper by Yu et al. (2024), we would not put much insight into the classification of the QPO nature. However, by the values of both $Q$-factor and RMS, it can be said that the QPO is type-C in nature for all the exposure IDs that have been presented here.
3.2 Spectral Properties
Along with studying the timing properties, studying the spectral properties gives a clearer idea about the nature of the source. With the available Insight-HXMT data, we analyzed the source for a total of 32 exposure IDs. These exposure IDs are marked with a ‘*’ sign in the 1st column of Table 2. This source hasn’t been extensively analyzed spectrally. We perform an extensive study of this source’s spectral properties for the first few days after the onset of its outburst. We started our spectral analysis on MJD 60181.4. We have used LE+ME+HE in the $2-150$ keV energy band (LE in 2–10, ME in 10–35, and HE in 35–150 keV) for our spectral fitting for all these exposure IDs.
Model Fitted Insight-HXMT Spectra
Refer to captionRefer to caption
Figure 5: Model fitted spectrally analyzed unfolded spectra. (a) is or observation ID. P061433800501 (Exposure: P061433800501-20230902-01-01) for which the gabs model was necessary to achieve the best fit. (b) is for the observation ID. P061433800601 (exposure: P061433800601-20230904-01-01) for which the gabs model was not needed to achieve the best fit.
We first fitted the data with phenomenological diskbb and power-law models. We used the multiplicative tbabs model to account for the interstellar absorption. Our model read as constant*tbabs(diskbb + powerlaw) in XSPEC. The constant is taken to normalize between three different instruments in LE, ME, and HE. We found that the fit was not statistically acceptable. There was a clear sign of the Gaussian feature in the unfolded spectrum around $5-6$ keV. Also, there was a hump-like feature $\sim 20$ keV. Besides these features, the absorption feature was noticed below 2 keV. Thus, we refitted using the model combination as constant*tbabs*gabs(diskbb + power-law + ga + pexrav). Here, the gabs model takes care of the low energy absorption feature, while the ga and pexrav models take care of the contribution from the Gaussian and hump-like nature, as mentioned above. In Fig. 5a, we show the best-fitted unfolded spectrum using this model combination for the observation ID P061433800501 (Exposure ID: P061433800501-20230902-01-01), where $\chi^{2}/DOF\sim 1$. After some exposure, we found that the gabs model was not needed anymore to achieve the best fit. Then, the model combination in XSPEC reads as constant*tbabs(diskbb + power-law + ga + pexrav). In Figure 5b, we show an unfolded best-fitted spectrum using the later model combination. This is for the observation ID. P061433800601 (exposure ID: P061433800601-20230904-01-01), for which $\chi^{2}/DOF\sim 0.9$.
Model Fitted Insight-HXMT Spectra
Refer to captionRefer to caption
Figure 6: Variation of spectrally fitted properties (a) hydrogen column density ($n_{H}$), (b) inner-disk temperature ($T_{in}$), (c) diskbb normalization, (d) Photon Index ($\Gamma$), (e) reflection fraction (Rref), and (f) cutoff energy (Ecut) with time.
[…]
3.3 Correlation between spectral and timing properties
Photon Index ($\Gamma$) vs QPO frequency ($\nu_{qpo}$)
Refer to captionRefer to captionRefer to caption
Figure 7: Variation of photon index of power-law ($\Gamma$) with QPO frequency ($\nu_{qpo}$) for (a) LE, (b) ME, and (c) HE instruments. Here, $\Gamma$ is the spectral fitted index using combined LE+ME+HE data.
The QPO frequency could be produced from the oscillation of the Compton corona, which could also be the origin of the hard power-law photons that extend the spectrum to very high energies. They could be correlated to each other. It has been previously seen that the QPO frequency ($\nu_{qpo}$) and the photon index of power-law ($\Gamma$) show a positive correlation for some of the sources during their active outbursting phase (Vignarca et al. 2003; Shaposhnikov & Titarchuk 2009; Stiele et al. 2013). We also tried to find out if there are any correlations between these two properties. In Figure 7, we show the variations of the photon index with the QPO frequency for the three energy bands. Considering these points, we have found a positive correlation between $\Gamma$ and $\nu_{qpo}$. Using both the Pearson-Linear and Spearman-Rank correlation methods, we found correlation coefficients for the three energy bands as below.
Pearson-Linear Spearman-Rank
LE $0.65\pm 0.15$ $0.65\pm 0.19$
ME $0.65\pm 0.15$ $0.65\pm 0.19$
HE $0.64\pm 0.15$ $0.64\pm 0.19$
4 Discussions
The Galactic black hole Swift J1727.8-1613 was monitored daily by various X-ray satellites like NICER, NuSTAR, Insight-HXMT, etc. Multi-wavelength follow-up observations were also carried out for this source. We have taken HXMT data for both our spectral and timing analysis from MJD 60181.4 to 60198.0. The evolution of the light curve makes the outburst nature of the source a fast-rise slow decay type. After the onset of the outburst, it rose to its peak very fast within 5 days. Using the $0.01~{}s$ time-binned light curves from the three instruments of HXMT (LE, ME, and HE), we studied the source’s timing properties by analyzing the light curves. Using the spectra files from these three instruments we also studied the spectral properties of this source by analyzing combined LE+ME+HE spectra in the $2-150$ keV broad energy range.
4.1 Propagation of Shock
Figure 8: Variation of shock properties with time in accordance with $\nu_{qpo}$. In (a) the variation of $\nu_{qpo}$ with time and in (b) and (c) we show the variations of the shock location ($X_{s}$), and shock speed ($v$) with time.
Quasi-periodic oscillation is one of the most important and common phenomena for stellar mass black holes. For this newly discovered source, we analyzed light curves of a total of 342 light curves (114 for LE, ME, and HE each). However, in some of the exposures, light curves were not produced properly. There were a total of 92 exposures which commonly produced LE, ME, and HE light curves. We list those data in Table 3. Our analysis is final for 276 data. We found the presence of QPO in all of these light curves and they show rapid evolution in frequency in this short analysis period. The QPO frequency ($\nu_{qpo}$) showed an increase in frequency within even a single day. This is one of the advantages of using Insight-HXMT as it provides multiple exposures for one observation ID which helps to study the short-scale variations of such important physical properties of the source. The evolution of the QPO frequency has been discussed broadly in the result section. It showed an increasing trend from the start of our analysis. From the variation of QPO frequency, (%)RMS, and $Q$-factor, the nature of the QPO is designated to be type-C. Although this is a well-established topic in the literature, the origin of QPO is still a matter of debate. Many models try to explain the type-C QPO phenomenon, e.g., the Lense-Thirring effect (Stella & Vietri 1998), accretion-ejection instability (Tagger & Pellat 1999), oscillation of the disk (Titarchuk & Osherovich 2000), the misalignment between the spin and binary rotational axis (Ingram et al. 2009, and references therein).
We would like to focus on a physical scenario that the QPOs could have originated from the shock instabilities in advective flows around black holes. Black hole accretion is transonic in which multiple sonic points could exist (Chakrabarti & Titarchuk 1995). Depending on the flow properties, there is the possibility of the formation of standing shocks (Singh et al. 2022 and references therein). The matter from the companion need not necessarily be only Keplerian. There could be a supply of matter which has angular momentum distribution deviated from the Keplerian one. This is the sub-Keplerian component. Such a lower angular momentum component of infalling matter gets accreted in a free-fall timescale. Due to the tug of war between the pressure of gravitational force and centrifugal force, this matter could virtually stop from falling onto the BH at some distance from the BH and undergo shock transition, creating a post-shock region. This approach could very well describe the spectral and temporal properties of BH, as it has been shown previously in both theoretical (Chakrabarti & Titarchuk 1995) and observational (Debnath et al. 2014; Mondal et al. 2014, Chatterjee et al. 2020, 2021, 2023) studies. The change in size of such post-shock region can describe the spectro-temporal properties during an outburst. According to Molteni et al. (1996), the oscillation in the shock region could have originated due to the presence of cooling. This was later shown in simulation works by Garain et al. (2014). The oscillation of the shock region could have originated due to the resonance condition between the cooling timescale due to the Comptonization process and the infall timescale in the coronal region (see Chakrabarti et al. 2015). Additionally, it has also been observed by the authors that once the QPO sets in, it gets locked for a week or more depending on the above condition. From the expressions of infall and cooling timescales, we get that the QPO frequency ($\nu_{qpo}$) could be given as (Molteni et al. 1996; Chakrabarti & Manickam 2000; Chakrabarti et al. 2004),
$\nu_{qpo}=\frac{c^{3}}{2GM_{BH}}\frac{1}{RX_{s}(X_{s}-1)^{1/2}}Hz,$ (1)
where $c$ and $G$ are the speed of light and gravitational constant respectively. $M_{BH}$ represents the mass of the black hole, $X_{s}$ is the distance of the shock, or called shock location. $X_{s}$ is in the units of Schwarzschild radius ($r_{s}=2GM_{BH}/c^{2}$). $R$ is the ratio of matter densities in post-shock to pre-shock regions ($\rho_{+}/\rho_{-}$) and is known as the compression ratio. From our timing analysis, we already extracted the information about QPO frequency ($\nu_{qpo}$). Using the above relation, we estimated the shock location during the outburst. According to Chakrabarti et al. (2005), $X_{s}$ could be located anywhere from 10 to 1000 $r_{s}$. The shock stays very far away $\sim 1000~{}r_{s}$ when the spectral nature of the ongoing outburst is hard and becomes smaller gradually in the process of cooling. Using the relation (1), we found that the shock was far at the start of the analysis period (Fig. 8b). It then rapidly decreased as the QPO frequency increased rapidly during the initial few days. Close to MJD 60185, when the $\nu_{qpo}$ stopped its rapid increment, and the shock location started to decrease very slowly. The values of the shock location along with QPO frequency are given in Table 3. As equation (1) suggests, as the shock was bigger at the start of the outburst, the QPO frequency was small. As it decreased rapidly due to cooling, the $\nu_{qpo}$ increased rapidly. Then when the cooling was less efficient, the size of the shock changed very slowly, although decreasing. It made the $\nu_{qpo}$ increase with time but not so rapidly as before.
If the time of the first exposure ID was taken to be 0, and the shock location as $X_{s0}$ and after time $t$, if the shock is at $X_{s}$, then the velocity of the movement of the shock would be given by the equation (Chakrabarti et al. 2005)
[…]
Using this relation, we estimated the velocity of shock. This is given in Fig. 8c. We find that the shock velocity has varied within this period. In the beginning, it attained higher values when the $X_{s}$ was also larger, however decreasing rapidly. It slowed down as the shock moved inwards, along with some fluctuations. This could be due to the limitation of assuming constant velocity of the movement of the shock.
4.2 Evolution of the physical configuration and spectral nature
From the spectral properties, we have seen that the $T_{in}$ was first low and it gradually increased. The exact opposite trend was observed for the Normalization of the diskbb model. The norm of diskbb model varies as
[…]
where, $R_{in}$, $D_{10}$ are the inner-disk radius of the accretion disk and the distance of the source in $10$ kpc unit. $i$ is the inclination angle of the accretion disk to the observer. Considering, $D_{10}$ and $\theta$ do not change during an outburst, the relation could be shortened like $R_{in}\sim N_{diskbb}^{1/2}$. Thus as the $N_{diskbb}$ decreased, it means the inner edge of the disk moved closer to the black hole. This is also supported by the fact that $T_{in}$ increased as the disk came closer.
This also supports the properties we showed from the timing analysis. The shock forms at a distance ($X_{s}$) when the sub-Keplerian matter is halted due to the counterbalance between the pressures of gravitational force and the centrifugal force. This forms the hot Compton cloud, as it becomes hot and puffed up in time. The hot Compton cloud will intercept some of those disk photons and there will be scattering in the process of inverse-Comptonization that will increase the energy of those intercepted photons and will emit them at higher energies. In this way, the Compton cloud will be cooled due to energy loss. Due to this cooling, the Compton cloud will gradually shrink in size. As it continues shrinking in size, the disk will continue to move inwards. As we have seen from the timing properties, the shock location was larger at the start of the outburst. The inner-disk radius was also higher as $R_{in}\sim N_{diskbb}^{1/2}$ and $N_{diskbb}$ was higher. Then cooling starts as the disk starts to approach inwards. In this way, the Compton cloud was losing energy by up-scattering and as the cooling continued, the $X_{s}$ and $R_{in}$ were both moving closer to the black hole.
Looking at the spectral parameters, and also the location of the shock, we designate that the source was already in an intermediate state at the start of our analysis period. As the outburst progressed, it was transitioning towards softer states.
From the values of the parameter $\cos(i)$ of the pexrav model, we find that the inclination of the disk was $>60^{\circ}$, which makes this source a high inclination binary system.
The hydrogen column density ($N_{H}$) showed a variation, which is common for Galactic black hole binaries unless the system is heavily obscured, which has recently been found for the newly discovered BH Swift J151857.0-572147 (Mondal et al. 2024). The average $N_{H}$ has an average value of $0.27_{-0.17}^{+0.08}\times 10^{22}$ cm-2.
5 Summary and Conclusions
We have studied the timing and spectral properties of the BHC Swift J1727.8-1613 during its recent outburst in 2023. We show the evolution of the light curve along with its hardness ratio for the entire duration of the outburst using archival MAXI/GSC data. Studying the entire evolution of the outburst was not the scope of this paper. Using the Insight-HXMT data, we chose the first ten observation IDs for our analysis. We use all the exposures from those observation IDs and selective exposures for timing and spectral analysis, respectively. Using $0.01~{}s$ time-binned light curves from all three instruments of HXMT, i.e., LE, ME, and HE, we studied the QPO properties by producing a power density spectrum. The QPO properties were extracted with the use of Lorentzian model. For spectral analysis, we use LE + ME + HE spectrum files in the broad $2-150$ keV energy band. We found that the models i) constant*tbabs*gabs*(diskbb + power-law + gaussian + pexrav) (for those for which gabs was required), ii) constant*tbabs*(diskbb + power-law + gaussian + pexrav) fits the spectra best. From our analysis, we conclude that:
(i) Quasi-periodic oscillations were present during the entire period of our analysis, which showed evolution in QPO frequency.
(ii) The $Q$-factor and $RMS$ help designate the nature of the QPOs as type-C.
(iii) QPOs originated from the shock instabilities that was formed by matter having an angular momentum distribution deviated from the Keplerian distribution. As the shock moved inwards, the QPO frequency was observed to increase.
(iv) With the shock moving inwards, the spectral nature of the source was transitioning towards the softer states, after it started in the intermediate state at the onset of the analysis period.
(v) The spectral and temporal properties correlate to each other which makes our claim stronger about the origin of QPOs.
(vi) This source is in a high inclination binary system.
(vii) The hydrogen column density varied in the range of $0.10-0.35\times 10^{22}$ cm-2, which is quite common for interstellar absorption of Galactic black holes.
6 Data Availability
This work has made use of public data from several satellite/instrument archives and has made use of software from the HEASARC, which is developed and monitored by the Astrophysics Science Division at NASA/GSFC and the High Energy Astrophysics Division of the Smithsonian Astrophysical Observatory. This work made use of the data from the Insight-HXMT mission, a project funded by the China National Space Administration (CNSA) and the Chinese Academy of Sciences (CAS).
7 Acknowledgements
KC acknowledges support from the SWIFAR postdoctoral fund of Yunnan University. SM acknowledges the Ramanujan Fellowship (# RJF/2020/000113) by SERB-DST, Govt. of India for this research. CBS is supported by the National Natural Science Foundation of China under grant no. 12073021.